OpenStudy (neo92):

Prove that dz/du-dz/dv = xdz/dx - ydz/dy where z=f(x,y) and x=e^u -e^-y and y=e^-u - e^v

OpenStudy (anonymous):

Whoops!

OpenStudy (anonymous):

Prove what that you are Neo92 ??

OpenStudy (neo92):

this is a partial differentiation problem

OpenStudy (neo92):

i need help bros..

OpenStudy (anonymous):

Did you mean:\[ x=e^u -e^{-v} \]??

OpenStudy (anonymous):

Did you try messing around with the chain rule?

OpenStudy (neo92):

yep

OpenStudy (anonymous):

\[ \frac{\partial z}{\partial u} = \frac{\partial z}{\partial x}\frac{\partial x}{\partial u} + \frac{\partial z}{\partial y}\frac{\partial y}{\partial u} \]

OpenStudy (anonymous):

Just use this fact to brute force it. Think you can do it @Neo92 ?

OpenStudy (neo92):

well.. i'll give it a try. Thnks @wio ..